OpenStudy (anonymous):

How can I find the graph of the quadratic function y = 3x^2 + 12x + 14

OpenStudy (blurbendy):

you start by putting in values for x start with 1 and see what that gives you for y after you plot a view x and y coordinates, you will see how the graph starts to take shape. Also, x can be negative if you want it to.

OpenStudy (blurbendy):

a few* not view

OpenStudy (anonymous):

I don't understand

OpenStudy (blurbendy):

so if you let x = 1, then: y = 3(1)^2 + 12(1) + 14 y = 3 + 12 + 14 y = 29 you get the coordinates (1,29) so plot them 

OpenStudy (blurbendy):

and you keep doing that for various x values